Exercise 1.1 from your textbook recommends that you try to make mistakes when experimenting with a new programming feature. This kind of experiment helps you remember what you read; it also helps when you are programming, because you get to know what the error messages mean. It is better to make mistakes now and on purpose than later and accidentally. (Downey, 2015, 7) For this Learning Journal, first answer the following questions based on Exercise 1.1. If you are trying to print a string, what happens if you leave out one of the quotation marks, or both? You can use a minus sign to make a negative number like -2. What happens for each of the following? >>> 2++2 >>> 2--2 >>> 2+-2 >>> 2-+2 In math notation, leading zeros are OK, as in 02. What happens if you try this in Python? What happens if you have two values with no operator between them? Next describe at least three additional Python experiments that you tried while learning Chapter 1. Show the Python inputs and their outputs, and explain what you learned from the results.
Answer #1

The answers are described in Python 3 as programming language:

1. a) Printing a string "Hello World" with quotation marks:

  • Input Code:

print("Hello World")

  • Output:

b) Printing a string "Hello World" without quotation marks:

If we put Hello World without one of the quotation marks or even both, the compiler will throw a syntax error as shown in the output. This is because if we put a word inside print function, the compiler processes it as a variable (example string in our case) and expects it to be declared.

  • Input Code:

print(Hello World)

  • Output:

  • Input Code:

print("Hello World)

  • Output:   

2. Evaluating mathematical expressions with two operators in between number:

  • Input Code:

print(2++2)

print(2--2)

print(2+-2)

print(2-+2)

  • Output:

​​​​​​​

This is because the compiler resolves its as:

  • 2+(+2)=2+2=4
  • 2-(-2)=2+2=4
  • 2+(-2)=2-2=0
  • 2-(+2)=2-2=0

3. Evaluating leading zeros in a number:

Python compiler treats a number with leading zero as invalid token.

  • Input Code:

print(401+004)

print(02)

  • Output:

​​​​​​​

4. If we have two values with no operator between them, the Python compiler treats it as a single value. An example is illustrated with Python code and output.

  • Input Code:

print(401+4)

print(4014)

  • Output:

​​​​​​​